OpenStudy (anonymous):

Can someone tell me how I set up this problem? What is the 7th term of the geometric sequence 4, -20, 100, ? -327,680 -393,216 -458,752 -524,288

OpenStudy (anonymous):

first term is 4 and the common ratio is \(\frac{-20}{4}=-5\) so the 7th term will be \[4\times (-5)^6\] and now you need a calculator

OpenStudy (anonymous):

okay let me pull it up on my laptop!!! :)) thank you!!!!!!!!

OpenStudy (anonymous):

62500

OpenStudy (anonymous):

wait I have a question - how did you get -20/4 = -5?

OpenStudy (anonymous):

i divided the second term by the first one

OpenStudy (anonymous):

oh that's right! that does equal five!! oh my gosh thank you so much it means so much!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw btw it equals(-5\)

OpenStudy (anonymous):

I'm going to try another problem on my own right now!!! :))

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

-5 is what i meant oops